Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.) μ = 5.3; σ = 1.8 P(7 ≤ x ≤ 9) =
Answer #1

Solution :

Given that ,

mean = = 5.3

standard deviation = = 1.8

P(7 ≤ x ≤ 9) = =P[(7 - 5.3) /1.8 < (x - ) / < (9 - 5.3) /1.8 )]

= P( 0.94< Z < 2.06)

= P(Z <2.06 ) - P(Z <0.94 )

Using z table,  

= 0.9803 - 0.8264

=0.1539